3. A. You buy seven copper futures at $2.79 per pound, where the contract size is 25,000 pounds. At contract maturity, copper is selling for $2.72 per pound. What is your profit or (loss) on the transaction?

B. You sell two aluminum futures at $2,332 per metric ton, where the contract size is 25 metric tons. At contract maturity, aluminum is selling for $2,315 per metric ton. What is your profit or (loss) on the transaction?

C. You buy 200 call option contracts on corn with a strike price of $3.60 per bushel at a quoted price of $0.07 per bushel, where the contract size is 5,000 bushels. At option expiration, corn sells for $3.65 per bushel. What is your net profit or (loss) on the transaction?

D. You buy 30 put option contracts on lean hogs with a strike price of $0.66 per pound at a quoted price of $0.05 per pound, where the contract size is 40,000 pounds. At option expiration, lean hogs sell for $0.59 per pound. What is your net profit or (loss) on the transaction?

E. You buy 30 put option contracts on lean hogs with a strike price of $0.05 per pound, where the contract size is 40,000 pounds. At option expiration, lean hogs sell for $0.68 per pound. What is your net profit or (loss) on the transaction?
Answer #1

A.

S0 = $2.79

S1 = $2.72

Contract size (z) = 25,000

Number of contracts (N) = 7

As we are long on the futures contract,

Profit or loss = (S1 - S0 ) * z * N

= (2.72- 2.79) * 25,000 * 7

= - $12,250

Hence a loss of $12,250

B.

S0 = $2,332

S1 = $2,315

Contract size (z) = 25

Number of contracts (N) = 2

As we are short on the futures contract,

Profit or loss = (S0 - S1) * z * N

= (2,332- 2,315) * 25 * 2

= 850

Hence a profit of $850

C.

Strike = $3.6

S1 = $3.65

Option premium (p)= $0.07

Contract size (z) = 5,000

Number of contracts (N) = 200

As the call option is in the money, we would exercise the option

Profit or loss = Cash flow from exercising the option - premium paid

= (S1 - Strike) * z * N - p * z * N

= (3.65 - 3.6) * 5,000 * 200 - 0.07 * 5,000 * 200

= -20,000

Hence a loss of $20,000

D.

Strike = $0.66

S1 = $0.59

Option premium (p)= $0.05

Contract size (z) = 40,000

Number of contracts (N) = 30

As the put option is in the money, we would exercise the option

Profit or loss = Cash flow from exercising the option - premium paid

= (Strike - S1) * z * N - p * z * N

= (0.66- 0.59) * 40,000 * 30 - 0.05 * 40,000 * 30

= 24,000

Hence a profit of $24,000

E.

As at the expiration the option is out of the money, we don't exercise our option. Hence we do not incur any loss or profit.